OpenStudy (anonymous):

I need help... PLEASE!!!!! Simplify the radical √1 / 5 Anwsers: 1 / √5 1 / 1.5 √5 / 5

OpenStudy (mertsj):

\[\sqrt{\frac{1}{5}}=\frac{\sqrt{1}}{\sqrt{5}}=\frac{1}{\sqrt{5}}\times\frac{\sqrt{5}}{\sqrt{5}}=\frac{\sqrt{5}}{5}\]

OpenStudy (anonymous):

Thank You. !

OpenStudy (mertsj):

yw